OpenStudy (anonymous):

if the radius of the circle is increased by 10% then the area is increased by?

OpenStudy (anonymous):

The relation between Area and radius is A= pie r^2

OpenStudy (mrnood):

Original area = pi r^2 New area = pi (r *1.1)^2 = pi r^2 *1.1^2 so divide the new by th eold to find the increase in area

OpenStudy (anonymous):

why its 1.1?

OpenStudy (anonymous):

the answer must be 2.1% 21% 121% none of these

OpenStudy (anonymous):

Why it is 1.1 instead of 0.1 @MrNood ? I think it should be 0.1.

OpenStudy (anonymous):

exactly it should be 0.1

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Not so sure about this.

OpenStudy (perl):

the above post is correct

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

The new radius is equal to the old radius + .1(of the old radius) so New Radius= r + .1r= 1.1 r

OpenStudy (perl):

(pi * (1.1*r)^2 ) / ( pi * r^2)

OpenStudy (anonymous):

ok whats the answer then?

OpenStudy (anonymous):

1.1^2 = 1.21

OpenStudy (anonymous):

New Radius= r + .1r= 1.1 r didnt get this step

OpenStudy (perl):

r + 10% * r

OpenStudy (anonymous):

why we plus r

OpenStudy (perl):

since you want to increase r

OpenStudy (anonymous):

ok

OpenStudy (perl):

if i have a 100 dollars, and i want to increase by 10% 100 + 10% * 100

OpenStudy (anonymous):

If you have 30 dollars and I give you 10 % 0f 30 =.1 (30)=3 dollars , the you will have 30+3=33= 1.1 (30)

OpenStudy (perl):

so in general, the formula is initial amount + (percentage increase) * initial amount

OpenStudy (anonymous):

ok

OpenStudy (perl):

so here we have r + 10%*r , or r + .10 * r

OpenStudy (anonymous):

yeh the answer is 121%?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

yes, the area is increased by 121 %

OpenStudy (perl):

alternatively, by a factor of 1.21

OpenStudy (anonymous):

Thanks @perl for the logic :)

OpenStudy (perl):

your welcome

OpenStudy (anonymous):

ty:)

OpenStudy (perl):

eliasaab also deserves credit

OpenStudy (anonymous):

yeh thanks @eliassaab

OpenStudy (perl):

@eliassaab good explanation

OpenStudy (anonymous):

Thanks @eliassaab :)